OpenStudy (anonymous):

Line n has an x-intercept of -4 negative four and a y-intercept of 5. What is the equation of a line parallel to line n that passes through the point (2, 3)?

OpenStudy (anonymous):

@RyleeC

OpenStudy (secret-ninja):

Choices?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

you know it must end with +5 since the y int. is 5

OpenStudy (anonymous):

the slope is -5/4

OpenStudy (anonymous):

so y = -5/4x + 5

OpenStudy (secret-ninja):

Does this help?

OpenStudy (secret-ninja):

-5x+4y=2 would be your answer.

OpenStudy (anonymous):

positive slope sorry

OpenStudy (anonymous):

answer is y = 5/4x + 5

OpenStudy (secret-ninja):

What are your choices?

OpenStudy (anonymous):

a)y=5/4x-1/2 b)y=5/4x+1/2 c)y=-5/4x-11/2 d)y=-5/4x+11/2

OpenStudy (secret-ninja):

I'm not sure... maybe c or d?

OpenStudy (anonymous):

its none

OpenStudy (anonymous):

if the y int is 5 then it must end in +5

OpenStudy (secret-ninja):

I think I would go with B